# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ Appendix E Results about atoms
Assume that $[G]$ is an atom. We proceed by induction on the number $n$ of cms’s of $G$. When $n=0$, the lemma is immediate. Assume the lemma is true when $n\leq k$, for any integer $k\geq 0$, and we will prove it when $n=k+1$. Let $C$ be any cms of $G$. Since $[G]$ is an induced minor and an atom, at most one $C$-component of $G$ contains some vertex that is not contained in $[C]^{-1}$. Let $H$ be any $C$-component of $G$ that contains $C$, which exists since $C$ is a minimal separator. If no $C$-component of $G$ contains some vertex that is not contained in $[C]^{-1}$, then each vertex of $G\setminus C$ is contained in $[C]^{-1}$, and hence $[H]^{-1}$. Otherwise, we can wlog choose $C$ and $H$ such that $H$ additionally contains some vertex that is not contained in $[C]^{-1}$. In this case, each vertex of $G\setminus H$ is contained in $[C]^{-1}$, and hence $[H]^{-1}$. Therefore, in either case, $[G]$ is an extension of some induced minor $[H]_{A}$ that is an atom.
Next, observe that $H$ has strictly fewer cms’s than $G$. Therefore, by the inductive hypothesis, $[H]_{A}$ is an extension of some induced minor $M$ of some atom of $H$. Finally, the lemma follows from the observation that $[G]$ is an extension of $M$. ∎
@@
Lemma 60 (How a cms partitions atoms and cms’s).
For any cms $C$ of a graph $G$, the set of all atoms of $G$ is equal to the disjoint union of the sets of all atoms of all $C$-components of $G$. Also, if no proper subgraph of any cms of $G$ is a cms of $G$, then set of all cms’s of $G$ is equal to the disjoint union of the set $\{C\}$ and the sets of all cms’s of all $C$-components of $G$.
@@
If some cms $E$ of $H$ is such that some connected component of $H\setminus E$ does not contain any vertex of $S\setminus E$, then clearly $E$ is a cs of $G$, and thus $G$ is not as atom. ∎
@@
We proceed by induction on the number $n$ of cms’s of $G$. When $n=0$, the lemma is immediate, and so the base case is true. For any integers $k\geq 0$ and $0\leq j\leq k$, assume the lemma is true when $n=j$, and we will prove it when $n=k+1$. Since $k+1>0$, $G$ has a cms $C$. Consider the set $\{H_{i}\}$ of $C$-components of $G$. Lemma 60 shows that each graph $H_{i}$ has strictly fewer cms’s than $G$, and any cms $E$ of $H_{i}$ is a cms of $G$. Hence, no proper subgraph of $E$ is a cms of $H_{i}$. Therefore, the inductive hypothesis applies to each atom graph $\mathcal{G}(H_{i})$.

### v2

@@ Appendix C Results about atoms
Assume that $[G]$ is an atom. We proceed by induction on the number $n$ of CMSs of $G$. When $n=0$, the lemma is immediate. Assume the lemma is true when $n\leq k$, for any integer $k\geq 0$, and we will prove it when $n=k+1$. Let $C$ be any CMS of $G$. Since $[G]$ is an induced minor and an atom, at most one $C$-component of $G$ contains some vertex that is not contained in $[C]^{-1}$. Let $H$ be any $C$-component of $G$ that contains $C$, which exists since $C$ is a minimal separator. If no $C$-component of $G$ contains some vertex that is not contained in $[C]^{-1}$, then each vertex of $G\setminus C$ is contained in $[C]^{-1}$, and hence $[H]^{-1}$. Otherwise, we can wlog choose $C$ and $H$ such that $H$ additionally contains some vertex that is not contained in $[C]^{-1}$. In this case, each vertex of $G\setminus H$ is contained in $[C]^{-1}$, and hence $[H]^{-1}$. Therefore, in either case, $[G]$ is an extension of some induced minor $[H]_{1}$ that is an atom.
Next, observe that $H$ has strictly fewer CMSs than $G$. Therefore, by the inductive hypothesis, $[H]_{1}$ is an extension of some induced minor $M$ of some atom of $H$. Finally, the lemma follows from the observation that $[G]$ is an extension of $M$. ∎
@@
Lemma 62 (How a CMS partitions atoms and CMSs).
For any CMS $C$ of a graph $G$, the set of all atoms of $G$ is equal to the disjoint union of the sets of all atoms of all $C$-components of $G$. Also, if no proper subgraph of any CMS of $G$ is a CMS of $G$, then set of all CMSs of $G$ is equal to the disjoint union of the set $\{C\}$ and the sets of all CMSs of all $C$-components of $G$.
@@
If some CMS $E$ of $H$ is such that some connected component of $H\setminus E$ does not contain any vertex of $S\setminus E$, then clearly $E$ is a clique separator of $G$, and thus $G$ is not as atom. ∎
@@
We proceed by induction on the number $n$ of CMSs of $G$. When $n=0$, the lemma is immediate, and so the base case is true. For any integers $k\geq 0$ and $0\leq j\leq k$, assume the lemma is true when $n=j$, and we will prove it when $n=k+1$. Since $k+1>0$, $G$ has a CMS $C$. Consider the set $\{H_{i}\}$ of $C$-components of $G$. Lemma 62 shows that each graph $H_{i}$ has strictly fewer CMSs than $G$, and any CMS $E$ of $H_{i}$ is a CMS of $G$. Hence, no proper subgraph of $E$ is a CMS of $H_{i}$. Therefore, the inductive hypothesis applies to each atom graph $\mathcal{G}(H_{i})$.